# String islands, discrete theta angles and the 6D $\mathcal{N}=(1,1)$ string landscape

### v1

@@ A.1.2 Two required lemmas
Proof: First note that by Theorem 1 the lattice $I_{n,4}$ admits a primitive embedding into $\Gamma_{n,n+8}\simeq\Gamma_{n,n}\oplus E_{8}$. By Theorem 2, the orthogonal complement of this primitive embedding is a lattice with signature $(0,4+n)$ with discriminant form $q^{\prime}=-q$ where $q$ is the discriminant form of $I_{n,4}$. Define $I_{4+n}$ as the associated lattice with bilinear form multiplied by $-1$; its discriminant form is $q$. Since $E_{8}(-1)$ and $\Gamma_{4,4}$ have trivial discriminant form, both the LHS and RHS of (A.1) have discriminant form $q$ hence they are locally isomorphic. Since the LHS has sublattice $\Gamma_{4,4}$, theorem 3 implies that the isomorphism is global. $\blacksquare$
@@ A.2.1 Cyclic case
To see that $g^{\prime}$ truly acts holomorphically, recall that $M$ acts trivially on the right-moving states of the worldsheet comprising the spacetime gravitini. The RHS of (A.10) is then accompanied by spacetime characters completing the second sum over $E_{8}(-1)$ to the 1-loop partition function of the antiholomorphic $E_{8}$ sCFT. This factorization of sums appears for all $g^{p}$ insertions, $p=1,...,N-1$, and since the partition function of the $E_{8}$ sCFT is modular-invariant, the partition function of the whole theory is holomorphically factorized.
@@ A.3 Extension to heterotic strings and possible degeneracies
At the points in moduli space where the Narain lattice is polarized as in (A.12), the worldsheet CFT is holomorphically factorized with the left-moving part corresponding to a lattice CFT based on $N_{I}$. Application of the allowed symmetries $M$ together with appropriate shifts $v$, in a systematic way, recovers the classification of chiral bosonic CFTs with central charge $c=24$ [17, 19]. This establishes the relationship between thes latter theories and the 6D $\mathcal{N}=(1,1)$ heterotic asymmetric orbifolds.
@@ Appendix B Computation of discrete theta angles
where $p$ satisfies $11<p<n-$ 1 and is coprime with $n$, and $a_{i}$ are the coefficients of the cyclotomic polynomial
@@
$\mathbb{Z}_{5}$ $\begin{pmatrix}0&0&0&-1\\ 1&0&0&-1\\ 0&1&0&-1\\ 0&0&1&-1\end{pmatrix}$ $\begin{array}[]{c}\frac{1}{5}(-u_{1}-u_{2}-u_{3}-u_{4},3u_{1}-2u_{2}-2u_{3}-2u_{4},\\ 2u_{1}+2u_{2}-3u_{3}-3u_{4},u_{1}+u_{2}+u_{3}-4u_{4})\\ \end{array}$

### v2

@@ A.1.2 Two required lemmas
Proof: First note that by Theorem 1 the lattice $I_{n,4}$ admits a primitive embedding into $\Gamma_{n,n+8}\simeq\Gamma_{n,n}\oplus E_{8}$. By Theorem 2, the orthogonal complement of this primitive embedding is a lattice with signature $(0,4+n)$ with discriminant form $q^{\prime}=-q$ where $q$ is the discriminant form of $I_{n,4}$. Define $I_{4+n}$ as the associated lattice with bilinear form multiplied by 1-1; its discriminant form is $q$. Since $E_{8}(-1)$ and $\Gamma_{4,4}$ have trivial discriminant form, both the LHS and RHS of (A.1) have discriminant form $q$ hence they are locally isomorphic. Since the LHS has sublattice $\Gamma_{4,4}$, theorem 3 implies that the isomorphism is global. $\blacksquare$
@@ A.2.1 Cyclic case
To see that $g^{\prime}$ truly acts holomorphically, recall that $M$ acts trivially on the right-moving states of the worldsheet comprising the spacetime gravitini. The RHS of (A.10) is then accompanied by spacetime characters completing the second sum over $E_{8}(-1)$ to the 1-loop partition function of the antiholomorphic $E_{8}$ sCFT. This factorization of sums appears for all $g^{p}$ insertions, p=1,,N1p=1,...,N-1, and since the partition function of the $E_{8}$ sCFT is modular-invariant, the partition function of the whole theory is holomorphically factorized.
@@ A.3 Extension to heterotic strings and possible degeneracies
At the points in moduli space where the Narain lattice is polarized as in (A.12), the worldsheet CFT is holomorphically factorized with the left-moving part corresponding to a lattice CFT based on $N_{I}$. Application of the allowed symmetries $M$ together with appropriate shifts $v$, in a systematic way, recovers the classification of chiral bosonic CFTs with central charge $c=24$ [35, 24]. This establishes the relationship between these latter theories and the 6D $\mathcal{N}=(1,1)$ heterotic asymmetric orbifolds.
@@ Appendix B Computation of discrete theta angles
where $p$ satisfies 1<p<n11<p<n-1 and is coprime with $n$, and $a_{i}$ are the coefficients of the cyclotomic polynomial
@@
[…]